OpenStudy (anonymous):

logbase2 7 1.a rational number 2.an irrational number 3.a prime number 4.an integer

OpenStudy (anonymous):

a

OpenStudy (anonymous):

b

OpenStudy (anonymous):

\[\frac{ \log7 }{ \log2 }=\frac{ 0.845 }{ 0.3010 }=2.80730879\]decide what it is

OpenStudy (anonymous):

lets do it this way log base 2 (7)=x 2^x=7 it may be rational or irrational considering the long nature of above result i think its irrational

OpenStudy (anonymous):

It is 'b' but do you understand why? logbase 2 of 7 means what power 'x' can you raise the number '2' to, and will equal 7? so\[2^{x}=7\] and you are solving for x now if you list the first few of the sequence you get \[2^{1}=2\]\[2^{2}=4\]\[2^{3}=8\] So already 8 is bigger than 7 So you know it can't be an integer, which also means it can't be a prime and I trust you are familiar with 2 to the power of a fraction? Meaning it can only be 'irrational'

OpenStudy (anonymous):

she could have guessed it from the final answer its b but still nice job sarah

OpenStudy (anonymous):

Assume that \(\bf log_2(7)= \frac{m}{n}\) for positive integers m and n. Then:\[\bf 2^{\frac{ m }{ n }}=7 \implies 2^m = 7^n\]This is clearly a contradiction because the left side is always even and the right side is always odd. Not only that, but if we look at the graph of \(\bf 2^x\) and \(\bf 7^x\), we notice that they only ever intersect when x = 0 which implies that \(\bf m=n=0\) but then \(\bf \frac{m}{n}=\frac{0}{0}\) which is not possible hence \(\bf 2^m=7^n\) is a contradiction. This proves that \(\bf log_2(7)\) is irrational by contradiction.

OpenStudy (anonymous):

@23ayesha1994

OpenStudy (anonymous):

thanks chandan and sara

OpenStudy (anonymous):

also genius

OpenStudy (anonymous):

@sarahusher You're method of arriving at the answer being irrational is inconclusive but irrational is still the correct answer :).

OpenStudy (anonymous):

@sarahusher i was taking the wrong ones out it was certainly not c or d hence i said so